# The CONDOR Observatory: A Gamma-Ray Observatory with a 100 GeV Threshold at 5300 Meters Above Sea Level

### v1

@@ Abstract
We present the design of the COmpact Network of Detectors with Orbital Range (CONDOR), a proposed high-altitude gamma-ray and cosmic-ray (CR) observatory set to become the highest of its kind. Planned for installation at Cerro Toco in the Atacama Desert, Chile, at 5300 meters above sea level (m.a.s.l.), CONDOR is optimized to operate in the 100 GeV to 1 TeV range using the extensive air-shower technique. The design prioritizes simplicity, modularity, and robustness to ensure reliable performance in a harsh environment. The CONDOR array has a fill factor of 90% and consists of 6000 plastic scintillator panels, each approximately 1 m2, read out by wavelength-shifting fibers and SiPMs. The readout electronics are based on flash ADCs, with White Rabbit technology ensuring time synchronization. We present an analysis of angular reconstruction and particle discrimination using CORSIKA-simulated CR showers, developing methods to reconstruct incident angles and distinguish between gamma-ray and proton CR events. CONDOR’s exceptional altitude and compact design enable it to meet the target 100 GeV threshold, complementing other ground-based observatories and overlapping with satellite detection ranges. CONDOR has the potential to support an extensive research program in astroparticle physics and multimessenger astronomy from the Southern Hemisphere, operating in all-sky mode 24/7.
@@ 2 The CONDOR Observatory
A possible layout of the WR network for CONDOR could be as follows. Assuming 6000 CONDOR units, about 600 White Rabbit nodes will be distributed across the CONDOR area through rugged outdoor optical fiber, distributing pulse-per-second (PPS) signals with sub-nanosecond accuracy to groups of about 10 panels. This will be achieved through a passive “fanout” board that distributes the PPS signal to each panel readout board via an SMA input, reducing the costs associated with additional WR nodes. A central DAQ hut will host a standard electronic rack that hosts necessary WR switches, which we estimate would be about 35, to synchronize 600 WR nodes. A master WR switch connects to an outdoor GPS antenna, although a rubidium atomic clock could serve as an alternative. The central hut will also host an industrial PC running Linux connected to the internet via Ethernet using the Atacama Astronomical Park network.
The power consumption of the CONDOR observatory is expected to be primarily driven by the WR nodes. Each WR node consumes less than 7 W, and each WR switch consumes less than 80 W. The DAQ industrial computer is expected to require approximately 250 W. In comparison, the power consumption of the CONDOR units is negligible (on the order of 1 mW per unit). Power will be supplied either by solar panels and batteries or through alternative solutions provided by the infrastructure of the Atacama Astronomical Park.
@@ 3 Expedition
We conducted an expedition and site-selection visit to the Atacama Astronomical Park in January 2023. After carefully evaluating potential sites based on criteria such as terrain suitability, logistical considerations, and other factors, we selected a location at Cerro Toco ($22^{\circ},57^{\prime},25^{\prime\prime},\text{S},;67^{\circ},46^{\prime},54^{\prime\prime},\text{W}$) as the proposed site for the CONDOR observatory (see Figure 5).
@@
Overall, the CONDOR site-selection expedition and chain test successfully completed its mission. This expedition marked an important first step toward demonstrating the remote operation of key components of the CONDOR under realistic conditions and finalizing the selection of the CONDOR observatory site. Additional results will be presented in a separate publication.
@@ 4.1 Monte Carlo Generator
4.2 Zenith angle reconstruction
We utilized the simulated data described in the previous section to reconstruct the zenith angle of the primary CR. Our approach involves fitting the shower front’s planar structure in the $(x,y,t)$ coordinate system, capturing the arrival times and positions of secondary particles detected across the array. This methodology allows us to determine the best-fit plane representing the shower front and infer the zenith angle.
@@ 4.2 Zenith angle reconstruction
$\theta=\arccos\left(\frac{\vec{n}\cdot\hat{z}}{|\vec{n}||\hat{z}|}\right)$
@@
This analysis was conducted for photon-initiated and proton-initiated showers, covering an energy range of 20 GeV to 800 GeV for photons and 150 GeV to 800 GeV for protons.
@@ 4.3.3 Efficiency, Purity, and Error Characterization
[…]
@@ 5 Summary and Discussion
Additionally, we implemented a Monte Carlo-based tagging system for gamma-ray and proton event discrimination. Our approach successfully classifies CR events with high efficiency and purity by employing statistical templates derived from simulated showers. The simulation results confirm that the observatory’s high-altitude location significantly enhances particle detection rates by reducing atmospheric attenuation. This advantage translates into a more precise measurement of CR properties, complementing both satellite-based and ground-based detection efforts.
@@
We envision building the CONDOR Observatory as the world’s highest-altitude gamma-ray and cosmic-ray observatory, distinguished by its exceptionally low energy threshold. Once completed, it will advance astroparticle physics and multimessenger astronomy [31] by measuring gamma rays in the 100 GeV–1 TeV range while operating continuously 24/7 in an all-sky mode. In addition, it will complement other EAS gamma-ray observatories that target higher energies at lower altitudes [7, 8, 10, 9] and Cherenkov imaging telescopes [32] that operate with narrower fields of view only at night.
@@ Acknowledgments
This work was funded by ANID PIA/APOYO AFB230003, Proyectos Internos de Investigación Multidisciplinarios USM 2024 PI__\__M__\__24__\__02 and Proyecto Milenio-ANID: ICN2019__\__044. The UC Riverside team work was supported by UC Riverside funding. We would like to thank the CLASS telescope crew for hosting us during our expedition and sharing their power and internet, the PAA personnel for assisting us during our visit, Miguel Mostafá for sharing his wisdom and enthusiasm, and Matt Kauer and Chris Wendt for their various forms of support for our design inspired in IceTop.

### v2

@@ Abstract
We present the design of the COmpact Network of Detectors with Orbital Range (CONDOR), a proposed high-altitude gamma-ray and cosmic-ray (CR) observatory set to become the highest of its kind. Planned for installation at Cerro Toco in the Atacama Desert, Chile, at 5300 meters above sea level (m.a.s.l.), CONDOR is optimized to operate in the 100 GeV to 1 TeV range using the extensive air-shower technique. The design prioritizes simplicity, modularity, and robustness to ensure reliable performance in a harsh environment. The CONDOR array has a fill factor of 90% and consists of 6000 plastic scintillator panels, each approximately 1 m2, read out by wavelength-shifting fibers and SiPMs. The readout electronics are based on flash ADCs, with White Rabbit technology ensuring time synchronization. We present an analysis of angular reconstruction and particle discrimination using CORSIKA-simulated CR showers, developing methods to reconstruct incident angles and distinguish between gamma-ray and proton CR events. CONDOR’s exceptional altitude and compact design enable it to meet the target 100 GeV threshold, complementing other ground-based observatories and overlapping with satellite detection ranges. CONDOR has the potential to support an extensive research program in astroparticle physics and multimessenger astronomy from the Southern Hemisphere, operating in all‐sky mode 24/7.
@@ 2 The CONDOR Observatory
A possible layout of the WR network for CONDOR could be as follows. Assuming 6000 CONDOR units, about 600 White Rabbit nodes will be distributed across the CONDOR area through rugged outdoor optical fiber, distributing pulse-per-second (PPS) signals with sub-nanosecond accuracy to groups of about 10 panels. To reduce system complexity and cost, we are investigating the use of a passive fanout board to distribute the PPS signal to each panel readout board via SMA connectors. While this approach offers simplicity, it may introduce timing dispersion, necessitating validation under controlled laboratory conditions. As part of future design iterations, we will also evaluate active fanout solutions and hierarchical switch-based synchronization architectures, guided by system-level performance requirements. A central DAQ hut will host a standard electronic rack that hosts necessary WR switches, which we estimate would be about 35, to synchronize 600 WR nodes. A master WR switch connects to an outdoor GPS antenna, although a rubidium atomic clock could serve as an alternative. The central hut will also host an industrial PC running Linux connected to the internet via Ethernet using the Atacama Astronomical Park network.
The power consumption of the CONDOR observatory is expected to be primarily driven by the WR nodes. Each WR node consumes less than 7 W, and each WR switch consumes less than 80 W. The DAQ industrial computer is expected to require approximately 250 W. Assuming the proposed node-to-detector mapping it yields to a total of approximately 7250 W. This corresponds to $\approx 12$ W per acquisition node or 1.2 W per detector unit on average. Power will be supplied either by solar panels and batteries or through alternative solutions provided by the infrastructure of the Atacama Astronomical Park.
@@ 3 Expedition
We conducted an expedition and site-selection visit to the Atacama Astronomical Park in January 2023. After carefully evaluating potential sites based on criteria such as terrain suitability, logistical considerations, and other factors, we selected a location at Cerro Toco ($22^{\circ},57^{\prime},25^{\prime\prime},\text{S};67^{\circ},46^{\prime},54^{\prime\prime},\text{W}$) as the proposed site for the CONDOR observatory (see Figure 5).
@@
The initial field test was limited to daytime, manual data collection, and was intended to validate basic DAQ operation under outdoor conditions. Environmental data (including temperature) were not yet integrated into the system. Future iterations will include onboard temperature sensors and support continuous, automated data collection, enabling precise temperature-rate correlations and dynamic bias adjustment.
@@
Additionally, we are developing a system to integrate LED-based calibration sources in each detector module. This will enable periodic in-situ characterization of SiPM gain, dark noise, and crosstalk, similar to the “sensor finger plot” approach described in [22]. These features are essential for optimizing the photon detection performance and ensuring long-term stability of the detector modules.
Overall, the CONDOR site-selection expedition and chain test successfully completed its mission. This expedition marked an important first step toward demonstrating the remote operation of key components of CONDOR under realistic conditions and finalizing the selection of the CONDOR observatory site. Additional results will be presented in a separate publication.
@@ 4.1 Monte Carlo Generator
4.2 Zenith angle and Energy reconstruction
We utilized the simulated data described in the previous section to reconstruct the zenith angle of the primary CR. Our approach involves fitting the shower front’s planar structure in the $(x,y,t)$ coordinate system, capturing the arrival times and positions of secondary particles detected across the array. This methodology allows us to determine the best-fit plane representing the shower front and infer the zenith angle. While conical shower front fits were also explored, they did not yield significant improvements in angular resolution at the low energies considered, due to the limited number of detected particles. As a result, planar fitting was adopted as the baseline reconstruction method.
@@ 4.2 Zenith angle and Energy reconstruction
[…]
@@
This analysis was conducted for photon-initiated and proton-initiated showers, covering an energy range of 20 to 800 GeV for photons and 150 to 800 GeV for protons.
The reconstructed energy values were obtained using a deep learning (DL) model designed to analyze data from simulated Extensive Air Showers (EAS) under the CONDOR Observatory conditions. Specifically, a hybrid architecture was employed that combines convolutional neural networks (CNNs) and transformers. This architecture merges the strengths of both models to overcome their individual limitations. CNNs are highly effective at extracting local features, such as the relationships between nearby detector hits, while transformers utilize a self-attention mechanism to capture global, long-range dependencies across the entire data sequence [29, 30]. The model processes the complete spatiotemporal sequence of detector activations, including the arrival time, deposited energy, and position of each hit, to extract complex patterns. This integrated approach leverages both the local and global information contained within the shower cascade.
The resolution of the energy reconstruction can be seen in Figure 11, where $\sigma(E)$ is the standard deviation of the reconstructed energy distribution at a given truth energy.
Figure 11: Energy reconstruction resolution as a function of true energy for four fixed incident energies, as obtained using a machine learning (ML) algorithm. The points are fitted with an exponential function to guide the trend. The resolution improves with increasing energy, as expected due to higher particle statistics and reduced relative fluctuations at larger energies.
@@ 4.3.3 Efficiency, Purity, and Error Characterization
Here, true selected represents the number of correctly tagged events, true total is the total number of true events, false selected is the number of misclassified events, and $\text{true selected}+\text{false selected}$ represents the total number of events tagged. These metrics provide a quantitative framework for evaluating the tagger’s performance in analyzing CR showers.
@@
Although the initial description focuses on proton-induced showers, the tagging algorithm was applied to both gamma and proton simulations. Figures 14 and 15 demonstrate the separation power between the two classes, showing that even with a simple template-based approach, stable discrimination is achievable across energy and zenith angle.
@@
4.4 CONDOR $\gamma$-ray sensitivity
Figure 16: Differential point-source sensitivity of the CONDOR $\gamma$-ray observatory for a one-year observation (90% duty cycle), compared with the sensitivities of HAWC, LHAASO, and CTA-South for equivalent exposure times. CONDOR shows enhanced sensitivity in the sub-TeV range ( 100–1000 GeV), complementing other observatories in this transitional energy domain.
Expected sensitivity of the CONDOR wide-field gamma-ray array for a one-year observation can be seen in Figure 16. The sensitivity was estimated by simulating the detector response to $\gamma$-ray flux, using a standard background-limited approach. An effective area of 5700 m² is assumed, with angular resolution and gamma/hadron discrimination performance taken from full detector simulations. The dominant background is assumed to be cosmic-ray protons, and their residual contamination is estimated using the ROC curve from Figure 15. The sensitivity calculation includes a realistic trigger condition, requiring at least 0.2 GeV of energy deposition in a single detection module, along with a coincidence of at least two adjacent stations within a 50 ns time window. This sequential trigger logic achieves a high detection efficiency (>95%) for gamma-induced showers, while significantly suppressing random background events. A duty cycle of 90% is assumed, consistent with continuous operation of ground-based air-shower arrays. In the same figure, we show for comparison the projected sensitivities of HAWC, LHAASO, and CTA-South for similar exposure times (1 year for HAWC and LHAASO; 50 hours for CTA-South as a standard reference). As expected for its design, CONDOR shows superior sensitivity in the sub-TeV energy range ( 100–1000 GeV), where its compact detector geometry and low-energy triggering allow efficient detection of low-energy air showers. In this regime, CONDOR complements existing instruments by filling the sensitivity gap between space-based telescopes (e.g., Fermi-LAT) and high-threshold ground-based observatories. At higher energies (>10 TeV), instruments like LHAASO and CTA outperform CONDOR due to their larger instrumented areas and optimized reconstruction for high-energy showers. Nevertheless, CONDOR maintains competitive sensitivity in the TeV range, demonstrating its potential for wide-field, continuous monitoring of transient and steady gamma-ray sources.
@@ 5 Summary and Discussion
Additionally, we implemented a Monte Carlo-based tagging system for gamma-ray and proton event discrimination. Our approach successfully classifies CR events with high efficiency and purity by employing statistical templates derived from simulated showers. The simulation results confirm that the observatory’s high-altitude location significantly enhances particle detection rates by reducing atmospheric attenuation. This advantage translates into a more precise measurement of CR properties, complementing both satellite-based and ground-based detection efforts. The current CONDOR design does not include muon tagging capabilities, which are known to enhance gamma/hadron separation as demonstrated by LHAASO and SWGO. Incorporating a muon-sensitive layer or underground veto system remains an important option for future detector upgrades.
@@
We envision building the CONDOR Observatory as the world’s highest‑altitude gamma‑ray and cosmic‑ray observatory, distinguished by its exceptionally low energy threshold. Once completed, it will advance astroparticle physics and multimessenger astronomy [34] by measuring gamma rays in the 100 GeV–1 TeV range while operating continuously 24/7 in an all‑sky mode. In addition, it will complement other EAS gamma-ray observatories that target higher energies at lower altitudes [7, 8, 10, 9] and Cherenkov imaging telescopes [35] that operate with narrower fields of view only at night.
@@ Acknowledgments
This work was funded by ANID PIA/APOYO AFB230003, Proyectos Internos de Investigación Multidisciplinarios USM 2024 PI_\_M_\_24_\_02 and Millennium Institute of Subatomic Physics at High Energy Frontier: ICN2019_\_044. The UC Riverside team work was supported by UC Riverside funding. We would like to thank the CLASS telescope crew for hosting us during our expedition and sharing their power and internet, the PAA personnel for assisting us during our visit, Miguel Mostafá for sharing his wisdom and enthusiasm, and Matt Kauer and Chris Wendt for their various forms of support for our design inspired in IceTop. Also the authors thank the High Performance Computing Cluster (HPCC) at DFIS, UTFSM, San Joaquín, Santiago, Chile